Let f(x)=|x-5| and g(x)=1/x
Write the given function as a composition function
using f and g.

H(x)=|(1/x)-5|
1
Expert's answer
2012-12-25T11:26:46-0500
H(x) = |(1/x)-5| = |g(x)-5|= f(g(x)) = f*g(x)